OpenStudy (anonymous):

lim for x-> ∞ from [(x+2)*e^(-1/x)] - x = ? wolfram says it is 1, but when we replace x with ∞ we have ∞ - ∞ , indeterminate form, so how should the expression be changed to avoid this?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} (x+2)\times e^{\frac{ -1 }{ x }} - x\] in mathematical terms :D

OpenStudy (shubhamsrg):

you may rewrite it as